# Constant-Size Cryptographic Evidence Structures for Regulated AI Workflows

### v1

@@ 5.4 Composition with TEEs
If events are generated inside TEEs [costan2016intel, sabt2015ted], the encoding functions $\phi_{i}$ can be defined to include attestation outputs or environment identifiers. For example, one field might be $H(\text{measurement}\parallel\text{config})$, where measurement is the TEE’s measurement of the loaded code. The abstraction provides a fixed-size container into which such identifiers can be hashed.
@@ 6.1 Implementation Overview
The prototype is implemented in and uses a standard 256-bit hash function and an Edwards-curve-based signature scheme. The implementation fixes a small constant $k$ (the number of fields) and a field length $\lambda$, but does not rely on any domain-specific interpretation of fields. Events are represented as structured records with randomly generated metadata; encoding functions $\{\phi_{i}\}$ extract and serialize appropriate components of each event before hashing.
@@ 6.2 Experimental Setup
Compiler Optimizing native compiler (release build)
@@ 6.5 Storage Considerations
7 Regulatory Alignment (High-Level)
Although we do not model any particular regulation in detail, constant-size evidence structures naturally support several common requirements:
@@ 7 Regulatory Alignment (High-Level)
Every security-relevant event is associated with a verifiable evidence item, providing a tamper-evident trail [hipaa].
@@
Fields that encode model identifiers, policy states, or configuration hashes support long-term tracking of model versions and configuration changes [muniswamy2006provenance, sultana2014secure].
@@ 8 Related Work
Schneier and Kelsey [schneier1999secure] and Bellare and Yee [bellare1997forward] proposed early designs for secure audit logs and forward-secure logging, focusing on cryptographic protection of log entries. Crosby and Wallach [crosby2009efficient] studied efficient data structures for tamper-evident logging. Our evidence structures build on these ideas but introduce a uniform, constant-size representation as a first-class abstraction tailored to regulated AI workflows.
@@
Merkle trees [merkle1980protocols] and time-stamping systems [haber1991time] underlie modern blockchain and transparency-log designs [nakamoto2008bitcoin, laurie2013certificate]. Our abstraction is designed to be ledger-friendly: evidence items are small and fixed-size, and our $\mathsf{Link}$ algorithm can produce succinct anchors suitable for external logs.
Ted execution environments.
TEEs such as Intel SGX [costan2016intel] and ARM TZone [sabt2015ted] provide hardware-isolated execution for security-critical code. Several systems use TEEs to protect data-processing pipelines and logging [hunt2018ryoan, ohrimenko2016oblivious]. Our design treats TEEs as optional components that can contribute identifiers or attestation digests to the encoding functions $\phi_{i}$, without changing the core evidence abstraction.
@@
Data provenance systems [muniswamy2006provenance, sultana2014secure] capture the lineage of data transformations, often at the file or database level. In regulated AI settings, provenance must be coupled with cryptographic integrity guarantees and regulator-aligned semantics. Constant-size evidence structures provide a compact, verifiable representation of key events and configurations that can coexist with richer provenance graphs.
@@
While secure logging systems [schneier1999secure, crosby2009efficient] and blockchain-based audits [laurie2013certificate, nakamoto2008bitcoin] provide integrity guarantees, they typically optimize for flexibility rather than uniformity. Variable-length records dominate existing frameworks, leading to:
@@ 9 Discussion and Limitations
  • Our security discussion is informal and follows established lines from secure logging and transparency systems; a full game-based formalization remains an interesting direction.
@@ 9.1 Lessons from Industrial Deployment
Retrofitting constant-size evidence structures into existing workflows can be challenging, particularly when legacy systems produce variable-length logs or lack well-defined event boundaries. In such cases, an intermediate translation layer may be needed to extract structured events and construct evidence items on-the-fly. This translation introduces a t boundary that must be carefully managed.
@@ 10 Conclusion
We believe that constant-size evidence structures can serve as a foundation for industry-wide standards in regulated AI audit trails. The abstraction’s emphasis on uniformity, cryptographic rigor, and hardware-friendliness makes it well-suited to emerging requirements in clinical trials, pharmaceutical manufacturing, medical AI, and financial compliance. We encourage practitioners, regulators, and standards bodies to consider fixed-size representations when designing next-generation compliance frameworks. Open discussion of these abstractions—including their security properties, performance trade-offs, and integration patterns—can accelerate the development of tworthy AI systems across healthcare, finance, and other high-stakes domains.
Future work includes formalizing security notions for evidence structures via game-based definitions, exploring richer encoding strategies that preserve privacy while supporting regulatory semantics (e.g., using zero-knowledge proofs or homomorphic commitments), and empirically studying how such abstractions integrate with full-scale AI systems in regulated environments. We also hope to engage with regulatory bodies and industry consortia to refine the design principles and to explore standardization opportunities.

### v2

@@ 6.4 Composition with TEEs
If events are generated inside TEEs [6, 21], the encoding functions $\phi_{i}$ can be defined to include attestation outputs or environment identifiers. For example, one field might be $H(\text{measurement}\parallel\text{config})$, where measurement is the TEE’s measurement of the loaded code. The abstraction provides a fixed-size container into which such identifiers can be hashed.
@@ 7.1 Implementation Overview
The prototype is implemented in Rust and uses a standard 256-bit hash function and an Edwards-curve-based signature scheme. The implementation fixes a small constant $k$ (the number of fields) and a field length $\lambda$, but does not rely on any domain-specific interpretation of fields. Events are represented as structured records with randomly generated metadata; encoding functions $\{\phi_{i}\}$ extract and serialize appropriate components of each event before hashing.
@@ 7.2 Experimental Setup
Compiler rustc (release mode, LTO enabled)
@@ 7.5 Storage Considerations
8 Regulatory Alignment
Although we do not model any particular regulation in detail, constant-size evidence structures naturally support requirements arising from frameworks such as HIPAA [27], the FDA’s electronic records rule [28], the EU AI Act [8], and financial reporting mandates [26]:
@@ 8 Regulatory Alignment
Every security-relevant event is associated with a verifiable evidence item, providing a tamper-evident trail [27].
@@
Fields that encode model identifiers, policy states, or configuration hashes support long-term tracking of model versions and configuration changes [16, 24].
@@ 9 Related Work
Schneier and Kelsey [22] and Bellare and Yee [4] proposed early designs for secure audit logs and forward-secure logging, focusing on cryptographic protection of log entries. Crosby and Wallach [7] studied efficient data structures for tamper-evident logging, and Ma and Tsudik [14] introduced new approaches using sequential aggregate signatures. Accorsi [1] surveys the landscape of secure logging protocols and identifies open challenges. Our evidence structures build on these ideas but introduce a uniform, constant-size representation as a first-class abstraction tailored to regulated AI workflows.
@@
Merkle trees [15], time-stamping systems [10, 2], and authenticated data structures [25] underlie modern blockchain and transparency-log designs [17, 13, 23]. Our abstraction is designed to be ledger-friendly: evidence items are small and fixed-size, and our $\mathsf{Link}$ algorithm can produce succinct anchors suitable for external logs.
Trusted execution environments.
TEEs such as Intel SGX [6] and ARM TrustZone [21] provide hardware-isolated execution for security-critical code. Several systems use TEEs to protect data-processing pipelines and logging [11, 19]. Our design treats TEEs as optional components that can contribute identifiers or attestation digests to the encoding functions $\phi_{i}$, without changing the core evidence abstraction.
@@
Data provenance systems [16, 24] capture the lineage of data transformations, often at the file or database level. In regulated AI settings, provenance must be coupled with cryptographic integrity guarantees and regulator-aligned semantics. Constant-size evidence structures provide a compact, verifiable representation of key events and configurations that can coexist with richer provenance graphs.
@@
While secure logging systems [22, 7] and blockchain-based audits [13, 17] provide integrity guarantees, they typically optimize for flexibility rather than uniformity. Variable-length records dominate existing frameworks, leading to:
@@ 10 Discussion and Limitations
  • Our security analysis assumes static key compromise; extensions to adaptive corruption models, concurrent multi-chain sessions, and composability frameworks remain open.
@@ 10.1 Lessons from Industrial Deployment
Retrofitting constant-size evidence structures into existing workflows can be challenging, particularly when legacy systems produce variable-length logs or lack well-defined event boundaries. In such cases, an intermediate translation layer may be needed to extract structured events and construct evidence items on-the-fly. This translation introduces a trust boundary that must be carefully managed.
@@ 11 Conclusion
We believe that constant-size evidence structures can serve as a foundation for industry-wide standards in regulated AI audit trails. The abstraction’s emphasis on uniformity, cryptographic rigor, and hardware-friendliness makes it well-suited to emerging requirements in clinical trials, pharmaceutical manufacturing, medical AI, and financial compliance. We encourage practitioners, regulators, and standards bodies to consider fixed-size representations when designing next-generation compliance frameworks. Open discussion of these abstractions—including their security properties, performance trade-offs, and integration patterns—can accelerate the development of trustworthy AI systems across healthcare, finance, and other high-stakes domains.
Future work includes extending the security model to adaptive adversaries who may corrupt signing keys during the protocol, analyzing composability of evidence structures under universal composability or simulation-based frameworks, and exploring richer encoding strategies that preserve privacy while supporting regulatory semantics (e.g., using zero-knowledge proofs or homomorphic commitments). We also plan to empirically study how such abstractions integrate with full-scale AI systems in regulated environments and to engage with regulatory bodies and industry consortia to refine the design principles and explore standardization opportunities.